Mickley Corporation produces two products, Alpha6s and Zeta7s, which pass through two operations, Sintering and Finishing. Each of the products uses two raw materials—X442 and Y661. The company uses a standard cost system, with the following standards for each product (on a per unit basis): Raw Material Standard Labor Time Product X442 Y661 Sintering Finishing Alpha6 2.0 kilos 3.0 liters 0.20 hours 0.90 hours Zeta7 5.0 kilos 4.0 liters 0.40 hours 1.00 hours Information relating to materials purchased and materials used in production during May follows: Material Purchases Purchase Cost Standard Price Used in Production X442 15,500 kilos $48,050 $2.90 per kilo 10,000 kilos Y661 16,500 liters $29,700 $1.90 per liter 14,500 liters The following additional information is available: The company recognizes price variances when materials are purchased. The standard labor rate is $22.00 per hour in Sintering and $20.50 per hour in Finishing. During May, 1,320 direct labor-hours were worked in Sintering at a total labor cost of $31,680, and 3,000 direct labor-hours were worked in Finishing at a total labor cost of $67,500. Production during May was 2,100 Alpha6s and 1,220 Zeta7s. Required:

1. Complete the standard cost card for each product, showing the standard cost of direct materials and direct labor.

2. Compute the materials price and quantity variances for each material.

3. Compute the labor rate and efficiency variances for each operation

Please Just help with what you can
Answer #1
Paart-1
Standard Standard Standard
Quantity or Hours Price or Rate Cost
Alpha6:
Direct materials-X442 2 kilos $2.90 per kilo $5.80
Direct materials-Y661 3 litres $1.90 per litre $5.70
Direct labor-Sintering 0.2 hours $22.00 per hour $4.40
Direct labor-Finishing 0.9 hours $20.50 per hour $18.45
tal $34.35
Zeta7:
Direct materials-X442 5 kilos $2.90 per kilo $14.50
Direct materials-Y661 4 litres $1.90 per litre $7.60
Direct labor-Sintering 0.4 hours $22.00 per hour $8.80
Direct labor-Finishing 1 hours $20.50 per hour $20.50
tal $51.40
Part-2
2. Material price variance = [SP-AP] x AQ purchased
Material X442 = [$2.90 - $3.10] x 15500 = $3100 U
Material Y661 = [$1.90 - $1.80] x 16500 = $1650 F
Material quantity variance = [SQ-AQ used] x SP
Material X442 = [{(2 x 2100) + (5 x 1220)} - 10000] x $2.90 = $870 F
Material Y661 = [{(3 x 2100) + (4.0 x 1220)} - 14500] x $1.90 = $6308 U
* X442 Actual Price per Kilos = 48050/15500=$3.10
**Y661, Actual Price Per Litre = 29700/16500=$1.80/ Litre
Direct material variance - Material X442
Material price variance 3100 U
Material quantity variance 1650 F
Direct material variance - Material Y661
Material price variance 870 F
Material quantity variance 6308UF
Part-3
3. Direct labor rate variance = [SR-AR] x AH worked
Sintering = [22 - 24] x 1320 = $2640 U
Finishing = [20.50 - 22.50] x 3000 = $6040 U
*Actual Rate/ Hour for Sintering :- $31680/1320=$24
**Actual Rate/ Hour for Finishing :- $67500/3000=$22.50
Direct labor efficiency variance = [SH-AH] x SR
Sintering = [{(0.20 x 2100) + (0.40 x 1200)} - 1320] x $22 = $9240 U
Finishing = [{(0.9 x 2100) + (1x 1200)} - 3000] x $20.50 = $1845 F
Direct labor variances - Sintering:
Labor rate variance 2640U
Labor efficiency variance 9240U
Direct labor variances - Finishing:
Labor rate variance 6040U
Labor efficiency variance 1845F